OpenStudy (anonymous):

Tina has to pay 4 percent of the purchase price toward closing costs for her mortgage. If the purchase price is $275,000, what is the closing cost amount?

OpenStudy (anonymous):

Okay, word problems can always be pesky, so let's get this into an equation. Any ideas on how to do that?

OpenStudy (anonymous):

4% of 275,000

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Sweet! Yep. Okay, any idea on how to figure out 4% of 275,000?

OpenStudy (anonymous):

4 times 275,000 ?

OpenStudy (anonymous):

Not quite. That would be a pretty big number. Let's put it into a ratio. \[\frac{ 4 }{ 100 } = \frac{ x }{ 275,000 }\]

OpenStudy (anonymous):

I meant multiply 4 by 275,000

OpenStudy (tamtoan):

4% = 4/100 so 4% of 275000 = (4/100) x 275000 = 4x2750 (since 275000/100 drop 2 zeros)

OpenStudy (anonymous):

and that would give you the percent right?

OpenStudy (anonymous):

Okay, so here 4% is represented as 4 over 100 (since percents are over one hundred.) So what you can do is do: 4 times 275,000 and take the answer and divide it by 100 (percent.)

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

110,000

OpenStudy (anonymous):

Not quite. You've got too many zeroes.

OpenStudy (anonymous):

11,000